OpenStudy (anonymous):

One number is five more than twice another. If their sum is decreased by ten, the result is twenty two . Find the numbers

OpenStudy (anonymous):

"twice another" 2y "five more than" +5 2y+5 = x There's one equation. "Their sum" x+y "is decreased by 10 -10 x+y-10 = 20 There's your second equation.

OpenStudy (raden):

the second, should be x+y-10 = 22

OpenStudy (anonymous):

Oh, right. Thank you.

OpenStudy (anonymous):

Let one number=x Other number=y Condition 1:One number is five more than twice another Twice the other number=2y Therefore, x = 2y + 5 (1st Equation) Condition 2:If their sum is decreased by ten, the result is twenty two Their sum=x + y Sum decreased by '10' = (x + y) - 10 Therefore, (x + y) - 10 = 22 (2nd Equation) Solving both the equations we get, x=23 y=9